OpenStudy (anonymous):

a box contains 6 blue pencils and 3 yellow pencils. what is the probability that the first 2 pencils randomly removed without replacement form the box will both be blue?

OpenStudy (anonymous):

6/9 times 5/8 = 5/12=41.7%